OpenStudy (joshoyen):

Solve by factoring. 4n^2 - 33n -5 = 0

OpenStudy (joshoyen):

im stuck on this :L

jimthompson5910 (jim_thompson5910):

notice how a = 4, b = -33 and c = -5 use the discriminant formula D = b^2 - 4ac to get D = b^2 - 4ac D = (-33)^2-4(4)(-5) D = 1169 Now take the square root of 1169 to get sqrt(1169) = 34.1906419945576 The square root result is NOT a whole number. So the discriminant isn't a perfect square when the discriminant isn't a perfect square, this means that we cannot factor

OpenStudy (joshoyen):

ohhh okay thanks!

jimthompson5910 (jim_thompson5910):

np